# Knowledge Graphs for the Life Sciences: Recent Developments, Challenges and Opportunities

### v1

@@ 5.1. What has been done: use cases and recent developments
Explainable AI for Knowledge Discovery. As introduced in Section 4, KGs can support knowledge discovery in life science, including the explainability of the process and the discovered units. In this view, Ritoski and Paulheim [150] explain that ontologies, linked data, and KGs are used in the interpretation step of a data mining process, e.g., for interpreting sequential patterns in patient data [79], or to describe subgroups in a semantic subgroup discovery process [172]. KGs can also serve both as the basis for knowledge discovery processes and the interpretation process. For example, Linked Open Data connecting drugs and adverse reactions can be analyzed with Hidden Conditional Random Fields to predict adverse drug reactions, where the paths from selected drugs to outcomes visually explain the prediction [89]. Similarly, Bresso et al. [13] leverage features extracted from KGs (interpretable features such as paths, neighbors, path patterns) and white box models (e.g., decision trees) to reproduce expert classifications of drugs causing or not specific adverse drug reactions. The rules extracted from the decision trees contain features that provide explanations for the molecular mechanisms behind these adverse reactions according to experts. Sousa et al. [162] employ KGs to explain both protein-protein interaction predictions and gene-disease association predictions based on shared semantic aspects.
Explainable AI for KG Construction The final use case considers the situation that XAI is applied to KGs themselves. We discussed the challenge to support human intervention in KG construction in Section 3.4. Recent KG construction gradually relies on data-driven, deep learning based methods to automatically induce knowledge from data. The deep learning models are opaque, and thus the process requires explainability. The resulting KG may not be accountable to be used for downstream applications. Trustworthy KG engineering is proposed in [194] to highlight the importance of embedding explainable AI and human intervention in the KG life cycle. XAI methods have been applied in many NLP related tasks (entity and relation extraction, entity resolution, link prediction, etc.) in KG construction from texts. The XAI methods rely either on feature-based explanations or knowledge-based explanations. While feature-based explanations try to infer explanations from the data or the models’ interpretation of the data, knowledge-based explanations aim to interpret the process with rules, reasoning paths, and structured contextual information. Rules and paths have mainly been used for explanation, especially for link prediction, a task comprehensively surveyed in [194].
5.2. What are the challenges?
@@
  • How to integrate KGs for better XAI, especially with recent deep learning and language model based methods? KG may provide better data provenance for the model output. This can ensure explainability for communicating the model to domain experts in data science applications [8]. In terms of recent generative LLMs, life science KGs, with careful curation based on scientific publications, may help to provide provenance data to the answers generated by LLMs. Studies need to understand to what extent, and how, LLMs can be applied to induce knowledge (e.g., by probing LLMs with biomedical ontologies [67]), which then may provide a foundation to create better approaches to integrate KGs with LLMs. Another area is neuro-symbolic methods which may provide models that are inherently more interpretable (see further discussions in Section 6.1). Also, regarding language models (especially LLMs), they are capable of generating fluent texts, which can potentially serve as textual explanation generators from symbolic knowledge for XAI. Meanwhile, a key issue is the hallucination of LLMs, and KGs may support better prompting, fine-tuning and interpretable inference of LLMs for higher decisiveness and trustfulness [137].
@@ 6. Discussion and Conclusion
6.1. Overall challenges and trends
Meanwhile, more scalable and efficient knowledge retrieval, query and reasoning systems, including life science KGs and mapping repositories, are still worthy of investigation and development.
Evolution and Quality Assurance of KGs. KGs need to be updated as new data and knowledge are emerging, and the schema and facts can easily become outdated or less useful for existing applications in life sciences. In terms of KG construction, we discussed ontology extension as a use case to address the evolution issue or emergence of new concepts and relations, and also instance matching to extend new instances for the KG. Updating KGs is also a prerequisite for life science knowledge discovery and knowledge discovery methods should be able to support the evolution of KGs with e.g., the capabilities of continuous learning and zero-shot learning. Quality assurance is another issue for KGs, including the tasks of knowledge error detection and correction, knowledge completion, knowledge canonicalization, etc. On the one hand, more effective KG quality assurance methods and systems should be developed, including schema and constraint languages for quality verification and learning-based models for prediction (e.g., [25] combines both for fact correction); on the other hand, knowledge discovery methods should be robust to noisy KGs by investigating e.g., robust KG embeddings and multi-modal representation learning.
Heterogeneity in KGs: Multi-domain and Multi-modality. KGs contain heterogeneous information, which brings challenges to their construction, representation, and reasoning. Different schema and data in KGs can have different focuses in their scopes and domains. Integrating data of different domains for building multi-domain KGs is difficult with challenges in e.g., ontology and data matching. Besides, recent studies have explored integrating different modalities to construct Multi-modal KGs [27, 125, 179], for instance text [136], images [182], etc. One challenge to address is how to learn effective machine learning models over multi-modal KGs fused from different sources (patients’ records, curated knowledge bases, and scientific literature) to support scientific discovery as well as KG construction and management. Another challenge is developing accurate and efficient knowledge representation approaches for texts and images in multi-modal KG construction. For example, careful consideration should be given to when to simply use an annotation property to associate an image with an entity, and when to use a property with specific semantics to connect an image and an entity.
Human Interaction and Explainability with KGs. In KG construction, human experts are required for many sub-tasks of KG construction and provide oversight [194]. In life science knowledge discovery, human experts are necessary to finally validate the predicted new knowledge. The whole process of interacting with KG in life sciences requires explainability, especially when sub-symbolic models (e.g., pre-trained language models) are used. How to generate clear explanations for human interaction and how to evaluate the quality of explanations remains a challenge, as well as how to achieve consensus regarding scientific understanding with automatically discovered knowledge when organizing knowledge in life science [132]. The recent growth of Neuro-Symbolic methods suggests their support for explainability [91, 92, 153]. A recent survey [92] summarizes XAI in bioinformatics with a chapter on knowledge-based explanations, whereas Karim [91, Chapter 8] provides a neuro-symbolic framework for KG construction and utilisation for medical experts’ decision making in the cancer domain. The approach presented in [153] is another recent example of neuro-symbolic integration for image classification with KG-based XAI in the cultural heritage domain.
Personalized and Customized KGs. A key challenge for KG construction is customisation, as we discussed in Section 3, to construct application-oriented KGs, where relevant sub-KGs have to be extracted for large-scale KGs (a.k.a. modualization) and integrated with other knowledge and data from different sources. Besides, many life science KGs are about individuals, e.g., patients in healthcare applications, where Personal Health KG enables the integration of instance-level (or patient-level) information and their computation is required [125]. An example is the Personal Health KG in [22] that supports the dietary recommendation for users, where the construction and population of the KG requires reusing and integrating existing ontologies, dietary guidelines, and time-series patient data. The other examples of KGs integrating patients’ EHR data [169, 16] are presented in Section 4.1. In personal KG construction, personal data should be protected. KG scalability should also be considered in order to be used on small devices such as cellphones. This is still a big challenge that has been rarely considered in using KGs in the life sciences.
Distributed KGs. The value of healthcare data for improving clinical knowledge and standard of care and the potential of semantic technologies to further enhance it are well recognized. However, a responsible use of healthcare data at the global level (beyond each healthcare provider and even each country) must take into account both legal and ethical issues in data sharing, privacy and security. Distributed knowledge graphs can mitigate these issues, by allowing for access control and privacy protection. Furthermore, distributed knowledge graphs can also address the challenges of scientific data ownership and stewardship by enabling the decentralized publishing of high quality data. Several approaches for federated querying and embedding of knowledge graphs have been proposed in recent years [26, 139, 160], however a wide adoption of semantic technologies in healthcare is still lacking, with a proliferation of terminological standards and a disconnection between data and meaning.
Representation Learning with KGs: Symbolic and Sub-symbolic Integration. Across the topics and use cases, we see the importance of transforming symbolic knowledge into sub-symbolic representations or combining both representations. The combination of both the neural and the traditional symbolic representation methods leads to a trend in neural-symbolic approaches in the field [12]. Recently, Pre-trained and Large Language Models provide new methods to transfer self-supervised learning from a vast amount of corpora to support KG construction, e.g., OntoGPT [18] and OntoLAMA [67]. LLMs are especially good at representing texts of life science publications in sub-symbolic spaces for semantic understanding. KGs may also provide a layer of explainability by validating the output of LLMs. A recent survey [137] proposes a roadmap for integrating LLMs and KGs. OntoProtein [195] is a recent example of how to integrate KGs into the process of pre-training LLMs in the bioinformatic domain, thus achieving improved results on protein-related knowledge discovery tasks. Also, geometry-informed representations of more formal KGs, especially in hyperbolic spaces or using complex geometric structures, e.g., [19, 102], can usually represent the structure of the KG with low dimensional vectors. Graph Neural Networks may also support the encoding of KG structures in a more explainable way with logical rules [31].
6.2. Conclusion
@@
Below we provide a list of key terms used in this paper, as well as their definitions and explanations. Note we mainly use the original sentences in the sources that are referenced as the definitions.
Description Logics: a family of knowledge representation languages that can be used to represent knowledge of an application domain. DLs differ from their predecessors, such as semantic networks and frames, in that they are equipped with logic-based semantics, the same semantics as that of classical first-order logic. Most ontologies are implemented in OWL, whose semantics are given by the Description Logic $\mathcal{SROIQ}$. [6]
TBox and ABox: the two components of domain knowledge in Description Logics, i.e., a terminological part called the TBox and an assertional part called the ABox, with the combination of a TBox and an ABox being called a knowledge base (KB). The TBox represents knowledge about the structure of the domain (similar to a database schema), while the ABox represents knowledge about a concrete situation (similar to a database instance). [6]
Semantic Networks: a graph structure for representing knowledge in patterns of interconnected nodes and arcs [163]. We use the term to denote a graph of concepts and relations without formal semantics.
Gene Ontology: The Gene Ontology (GO) knowledgebase provides a comprehensive, structured, computer-accessible representation of gene function, for genes from any cellular organism or virus [5, 29].
SNOMED-CT: Systematized Nomenclature of Medicine Clinical Terms (SNOMED CT) is a structured clinical vocabulary. It has a general and comprehensive coverage of clinical terms to support electronic healthcare systems and clinical applications. [40, 28]
UMLS (UMLS Metahesaurus and UMLS Semantic Networks): Unified Medical Language System (UMLS) is a repository of biomedical vocabularies developed by the US National Library of Medicine. The UMLS is composed of three “knowledge sources”, a Metathesaurus, a semantic network, and a lexicon. The UMLS Metathesaurus is a comprehensive effort for integrating biomedical ontologies through mappings. The UMLS Semantic Networks define the types or categories, or Semantic Types, of all Metathesaurus concepts and their relationships, or Semantic Relations. [10, 28]
ChEBI: Chemical Entities of Biological Interest (ChEBI) is a database and ontology containing information about chemical entities of biological interest. [64]
Symbolic vs. subsymbolic representations: Rooted in cognitive science, symbolic systems of human cognition are related to the representation and manipulation of symbols; sub-symbolic or connectionist systems are most generally associated with the metaphor of a neuron, e.g., perceptrons as an early system [97]. In terms of AI, symbolic systems contain logic-based and knowledge representations, while subsymbolic systems typically contain neural networks and deep learning based methods [43]. Neural language models and pre-trained language models [88] are also classified under subsymbolic systems.
Pre-trained and Large Language Models: Neural language modelling is the task of using neural network approaches to predict words from prior their contexts in a sequence. Pre-training is the process of learning some sort of representation (usually neural embedding based) of meaning for words or sentences by processing very large amounts of text (or other data in a sequence form, e.g., proteins and KG facts). This results in pre-trained language models. The dominating architecture for neural language modeling is Transformer-based models, including BERT, its domain specific versions, and later large variants, like the GPT series. The pre-trained language models of very large sizes are recently coined Large Language Models (LLMs). [88]
Neuro-symbolic representations: refers to the integration of neural networks and symbolic representations to design AI models that base their prediction on both data and knowledge. [43]
@@
Acknowledgements
We would like to thank Uli Sattler (University of Manchester) for proposing the topic of this paper and Terry Payne (University of Liverpool) for the useful comments on a previous draft. We would also like to thank the TGDK editors in chief for organizing this inaugural issue.

### v4

@@ 5.1 What has been done: use cases and recent developments
Explainable AI for Healthcare Practice. The utilization of AI in healthcare practice raises the concern of leaving life-critical decisions to black-box models [143, 167]. For example, in the field of precision medicine which aims at tailoring drug treatments and dosages to each patient, clinicians require more information from a model than a simple binary decision [7]. The interpretability and explainability of AI models is thus an essential characteristic to make outputs understandable and transparent. This would enforce both clinicians’ and patients’ trust in models by complementing (and not substituting) clinicians’ explanations [20, 143, 167]. To illustrate, this direction has been envisioned for several healthcare scenarios. Explainable AI models could support the experts in finding clinical trials that are appropriate based on patient history [167]. Counterintuitive or unreliable predictions that could have serious consequences could be explained, and thus prevented [167, 14, 90]. Some also envision such models to be used to explain and debunk healthcare-related misinformation [143]. As aforementioned, it is noteworthy that different kinds of explanations should be employed depending on the target audience, e.g., scientific explanations for evidence or trace-based explanations for treatment [20]. Explainable AI for Knowledge Discovery. As introduced in Section 4, KGs can support knowledge discovery in life science, including the explainability of the process and the discovered units. In this view, Ritoski and Paulheim [147] explain that ontologies, linked data, and KGs are used in the interpretation step of a data mining process, e.g., for interpreting sequential patterns in patient data [77], or to describe subgroups in a semantic subgroup discovery process [169]. KGs can also serve both as the basis for knowledge discovery processes and the interpretation process. For example, Linked Open Data connecting drugs and adverse reactions can be analyzed with Hidden Conditional Random Fields to predict adverse drug reactions, where the paths from selected drugs to outcomes visually explain the prediction [87]. Similarly, Bresso et al. [12] leverage features extracted from KGs (interpretable features such as paths, neighbors, path patterns) and white box models (e.g., decision trees) to reproduce expert classifications of drugs causing or not specific adverse drug reactions. The rules extracted from the decision trees contain features that provide explanations for the molecular mechanisms behind these adverse reactions according to experts. Sousa et al. [159] employ KGs to explain both protein-protein interaction predictions and gene-disease association predictions based on shared semantic aspects. Explainable AI for KG Construction The final use case considers the situation that XAI is applied to KGs themselves. We discussed the challenge to support human intervention in KG construction in Section 3.4. Recent KG construction gradually relies on data-driven, deep learning based methods to automatically induce knowledge from data. The deep learning models are opaque, and thus the process requires explainability. The resulting KG may not be accountable to be used for downstream applications. Trustworthy KG engineering is proposed in [191] to highlight the importance of embedding explainable AI and human intervention in the KG life cycle. XAI methods have been applied in many NLP related tasks (entity and relation extraction, entity resolution, link prediction, etc.) in KG construction from texts. The XAI methods rely either on feature-based explanations or knowledge-based explanations. While feature-based explanations try to infer explanations from the data or the models’ interpretation of the data, knowledge-based explanations aim to interpret the process with rules, reasoning paths, and structured contextual information. Rules and paths have mainly been used for explanation, especially for link prediction, a task comprehensively surveyed in [191].
5.2 What are the challenges?
@@
  • How to integrate KGs for better XAI, especially with recent deep learning and language model based methods? KG may provide better data provenance for the model output. This can ensure explainability for communicating the model to domain experts in data science applications [7]. In terms of recent generative LLMs, life science KGs, with careful curation based on scientific publications, may help to provide provenance data to the answers generated by LLMs. Studies need to understand to what extent, and how, LLMs can be applied to induce knowledge (e.g., by probing LLMs with biomedical ontologies [65]), which then may provide a foundation to create better approaches to integrate KGs with LLMs. Another area is neuro-symbolic methods which may provide models that are inherently more interpretable (see further discussions in Section 6.1). Also, regarding language models (especially LLMs), they are capable of generating fluent texts, which can potentially serve as textual explanation generators from symbolic knowledge for XAI. Meanwhile, a key issue is the hallucination of LLMs, and KGs may support better prompting, fine-tuning and interpretable inference of LLMs for higher decisiveness and trustworthiness [134].
@@ 6 Discussion and Conclusion
6.1 Overall challenges and trends
Meanwhile, more scalable and efficient knowledge retrieval, query and reasoning systems, including life science KGs and mapping repositories, are still worthy of investigation and development. Evolution and Quality Assurance of KGs. KGs need to be updated as new data and knowledge are emerging, and the schema and facts can easily become outdated or less useful for existing applications in life sciences. In terms of KG construction, we discussed ontology extension as a use case to address the evolution issue or emergence of new concepts and relations, and also instance matching to extend new instances for the KG. Updating KGs is also a prerequisite for life science knowledge discovery and knowledge discovery methods should be able to support the evolution of KGs with e.g., the capabilities of continuous learning and zero-shot learning. Quality assurance is another issue for KGs, including the tasks of knowledge error detection and correction, knowledge completion, knowledge canonicalization, etc. On the one hand, more effective KG quality assurance methods and systems should be developed, including schema and constraint languages for quality verification and learning-based models for prediction (e.g., [24] combines both for fact correction); on the other hand, knowledge discovery methods should be robust to noisy KGs by investigating e.g., robust KG embeddings and multi-modal representation learning. Heterogeneity in KGs: Multi-domain and Multi-modality. KGs contain heterogeneous information, which brings challenges to their construction, representation, and reasoning. Different schema and data in KGs can have different focuses in their scopes and domains. Integrating data of different domains for building multi-domain KGs is difficult with challenges in e.g., ontology and data matching. Besides, recent studies have explored integrating different modalities to construct Multi-modal KGs [26, 122, 176], for instance text [133], images [179], etc. One challenge to address is how to learn effective machine learning models over multi-modal KGs fused from different sources (patients’ records, curated knowledge bases, and scientific literature) to support scientific discovery as well as KG construction and management. Another challenge is developing accurate and efficient knowledge representation approaches for texts and images in multi-modal KG construction. For example, careful consideration should be given to when to simply use an annotation property to associate an image with an entity, and when to use a property with specific semantics to connect an image and an entity. Human Interaction and Explainability with KGs. In KG construction, human experts are required for many sub-tasks of KG construction and provide oversight [191]. In life science knowledge discovery, human experts are necessary to finally validate the predicted new knowledge. The whole process of interacting with KG in life sciences requires explainability, especially when sub-symbolic models (e.g., pre-trained language models) are used. How to generate clear explanations for human interaction and how to evaluate the quality of explanations remains a challenge, as well as how to achieve consensus regarding scientific understanding with automatically discovered knowledge when organizing knowledge in life science [129]. The recent growth of Neuro-Symbolic methods suggests their support for explainability [89, 90, 150]. A recent survey [90] summarizes XAI in bioinformatics with a chapter on knowledge-based explanations, whereas Karim [89, Chapter 8] provides a neuro-symbolic framework for KG construction and utilization for medical experts’ decision making in the cancer domain. The approach presented in [150] is another recent example of neuro-symbolic integration for image classification with KG-based XAI in the cultural heritage domain. Personalized and Customized KGs. A key challenge for KG construction is customization, as we discussed in Section 3, to construct application-oriented KGs, where relevant sub-KGs have to be extracted for large-scale KGs (a.k.a. modularization) and integrated with other knowledge and data from different sources. Besides, many life science KGs are about individuals, e.g., patients in healthcare applications, where Personal Health KG enables the integration of instance-level (or patient-level) information and their computation is required [122]. An example is the Personal Health KG in [21] that supports the dietary recommendation for users, where the construction and population of the KG requires reusing and integrating existing ontologies, dietary guidelines, and time-series patient data. The other examples of KGs integrating patients’ EHR data [166, 15] are presented in Section 4.1. In personal KG construction, personal data should be protected. KG scalability should also be considered in order to be used on small devices such as cellphones. This is still a big challenge that has been rarely considered in using KGs in the life sciences. Distributed KGs. The value of healthcare data for improving clinical knowledge and standard of care and the potential of semantic technologies to further enhance it are well recognized. However, a responsible use of healthcare data at the global level (beyond each healthcare provider and even each country) must take into account both legal and ethical issues in data sharing, privacy and security. Distributed knowledge graphs can mitigate these issues, by allowing for access control and privacy protection. Furthermore, distributed knowledge graphs can also address the challenges of scientific data ownership and stewardship by enabling the decentralized publishing of high quality data. Several approaches for federated querying and embedding of knowledge graphs have been proposed in recent years [25, 136, 157], however a wide adoption of semantic technologies in healthcare is still lacking, with a proliferation of terminological standards and a disconnection between data and meaning. Representation Learning with KGs: Symbolic and Sub-symbolic Integration. Across the topics and use cases, we see the importance of transforming symbolic knowledge into sub-symbolic representations or combining both representations. The combination of both the neural and the traditional symbolic representation methods leads to a trend in neural-symbolic approaches in the field [11]. Recently, Pre-trained and Large Language Models provide new methods to transfer self-supervised learning from a vast amount of corpora to support KG construction, e.g., OntoGPT [17] and OntoLAMA [65]. LLMs are especially good at representing texts of life science publications in sub-symbolic spaces for semantic understanding. KGs may also provide a layer of explainability by validating the output of LLMs. A recent survey [134] proposes a roadmap for integrating LLMs and KGs. OntoProtein [192] is a recent example of how to integrate KGs into the process of pre-training LLMs in the bioinformatic domain, thus achieving improved results on protein-related knowledge discovery tasks. Also, geometry-informed representations of more formal KGs, especially in hyperbolic spaces or using complex geometric structures, e.g., [18, 99], can usually represent the structure of the KG with low dimensional vectors. Graph Neural Networks may also support the encoding of KG structures in a more explainable way with logical rules [30].
6.2 Conclusion
@@
Below we provide a list of key terms used in this paper, as well as their definitions and explanations. Note we mainly use the original sentences in the sources that are referenced as the definitions. Description Logics: a family of knowledge representation languages that can be used to represent knowledge of an application domain. DLs differ from their predecessors, such as semantic networks and frames, in that they are equipped with logic-based semantics, the same semantics as that of classical first-order logic. Most ontologies are implemented in OWL, whose semantics are given by the Description Logic $\mathcal{SROIQ}$. [5] TBox and ABox: the two components of domain knowledge in Description Logics, i.e., a terminological part called the TBox and an assertional part called the ABox, with the combination of a TBox and an ABox being called a knowledge base (KB). The TBox represents knowledge about the structure of the domain (similar to a database schema), while the ABox represents knowledge about a concrete situation (similar to a database instance). [5] Semantic Networks: a graph structure for representing knowledge in patterns of interconnected nodes and arcs [160]. We use the term to denote a graph of concepts and relations without formal semantics. Gene Ontology: The Gene Ontology (GO) knowledgebase provides a comprehensive, structured, computer-accessible representation of gene function, for genes from any cellular organism or virus [4, 28]. SNOMED-CT: Systematized Nomenclature of Medicine Clinical Terms (SNOMED CT) is a structured clinical vocabulary. It has a general and comprehensive coverage of clinical terms to support electronic healthcare systems and clinical applications. [38, 27] UMLS (UMLS Metahesaurus and UMLS Semantic Networks): Unified Medical Language System (UMLS) is a repository of biomedical vocabularies developed by the US National Library of Medicine. The UMLS is composed of three “knowledge sources”, a Metathesaurus, a semantic network, and a lexicon. The UMLS Metathesaurus is a comprehensive effort for integrating biomedical ontologies through mappings. The UMLS Semantic Networks define the types or categories, or Semantic Types, of all Metathesaurus concepts and their relationships, or Semantic Relations. [9, 27] ChEBI: Chemical Entities of Biological Interest (ChEBI) is a database and ontology containing information about chemical entities of biological interest. [62] Symbolic vs. subsymbolic representations: Rooted in cognitive science, symbolic systems of human cognition are related to the representation and manipulation of symbols; sub-symbolic or connectionist systems are most generally associated with the metaphor of a neuron, e.g., perceptrons as an early system [94]. In terms of AI, symbolic systems contain logic-based and knowledge representations, while subsymbolic systems typically contain neural networks and deep learning based methods [41]. Neural language models and pre-trained language models [86] are also classified under subsymbolic systems. Pre-trained and Large Language Models: Neural language modeling is the task of using neural network approaches to predict words from prior their contexts in a sequence. Pre-training is the process of learning some sort of representation (usually neural embedding based) of meaning for words or sentences by processing very large amounts of text (or other data in a sequence form, e.g., proteins and KG facts). This results in pre-trained language models. The dominating architecture for neural language modeling is Transformer-based models, including BERT, its domain specific versions, and later large variants, like the GPT series. The pre-trained language models of very large sizes are recently coined Large Language Models (LLMs). [86] Neuro-symbolic representations: refers to the integration of neural networks and symbolic representations to design AI models that base their prediction on both data and knowledge. [41]